OpenStudy (anonymous):

Which function rule can be used to represent the area of a triangle with a base b 8 in. longer than twice the height h in the terms of the height? Choices will be posted below...

OpenStudy (anonymous):

\[A. A = 1/2bh\] \[B. A=1/2h(h+8)\] \[C. A=h^2+4h\] \[D. A=1/2(2h)(h+8)\]

OpenStudy (anonymous):

The answer is not A. It just isn't . . . it doesn't go with the problem.

OpenStudy (anonymous):

The answer is not C because it has nothing to do with multiplying the height by 4 and squaring it

OpenStudy (anonymous):

Because you are looking at twice the height, you have to multiply h by 2 The answer is B Does that make sense?

OpenStudy (anonymous):

oh yes ! thanks :D